OpenStudy (anonymous):

Solve the equation log10x = 1/4log10*16 +1/2log10*49 the log is base 10.

OpenStudy (anonymous):

\[\log(10x)=\log(10*2^4)^{\frac{1}{4}}+\log(10*7^2)^{\frac{1}{2}}\]

OpenStudy (anonymous):

\[\log(10x)=\log(10*2)+(\log10*7)\]

OpenStudy (anonymous):

\[\log(10x)=\log(20*70)\]

OpenStudy (anonymous):

I dont understand. I thought 10 was the base. Are we supposed to multiply the base?

OpenStudy (anonymous):

log10x is base 1/4log10*16 all 10 is base?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[\log _{10}(n)+\log _{10}(m)=(n*m)\]

OpenStudy (anonymous):

Oh I see. I understand now. Thank you.

OpenStudy (anonymous):

\[\log(10x)=\log(1400)\]

OpenStudy (anonymous):

10x=1400 \[x=\frac{1400}{10}=140\] x=140

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

I was wrong first line:\[\log(10x)=\log(10*(2^4)^\frac{1}{4}) + \log (10*(7^2)^{\frac{1}{2}})\]

OpenStudy (anonymous):

So what is the answer now?

OpenStudy (anonymous):

The same answer, I just rewite the first line that all